# Mixed-Type Tabular Data Synthesis with Score-based Diffusion in Latent Space

### v1

@@ Abstract
Recent advances in tabular data generation have greatly enhanced synthetic data quality. However, extending diffusion models to tabular data is challenging due to the intricately varied distributions and a blend of data types of tabular data. This paper introduces TabSyn, a methodology that synthesizes tabular data by leveraging a diffusion model within a variational autoencoder (VAE) crafted latent space. The key advantages of the proposed TabSyn include (1) Generality: the ability to handle a broad spectrum of data types by converting them into a single unified space and explicitly capture inter-column relations, (2) Quality: optimizing the distribution of latent embeddings to enhance the subsequent training of diffusion models, which helps generate high-quality synthetic data, (3) Speed: much fewer number of reverse steps and faster synthesis speed than existing diffusion-based methods. Extensive experiments on six datasets with five metrics demonstrate that TabSyn outperforms existing methods. Specifically, it reduces the error rates by 86% and 67% for column-wise distribution and pair-wise column correlation estimations compared with the most competitive baselines. Code has been made available at https://github.com/amazon-science/tabsyn.
@@ 1 Introduction
Figure 1: TabSyn wins: Our TabSyn method consistently outperforms SOTA tabular data generation methods across five data quality metrics.
@@
This paper presents TabSyn, a principled approach for tabular data synthesis that involves training a diffusion model in the latent space. TabSyn first transforms raw tabular data into a continuous embedding space, where well-developed diffusion models with Gaussian noises become feasible. Subsequently, we learn a score-based diffusion model in the embedding space to capture the distribution of latent embeddings. To learn an informative, smoothed latent space while maintaining the decoder’s reconstruction ability, we specifically designed a Variational AutoEncoder (VAE (Kingma & Welling, 2013)) model for tabular-structured data. Our proposed VAE model includes 1) Transformer-architecture encoders and decoders for modeling inter-column relationships and obtaining token-level representations, facilitating token-level tasks. 2) Adaptive loss weighting to dynamically adjust the reconstruction loss weights and KL-divergence weights, allowing the model to improve reconstruction performance gradually while maintaining a regularized embedding space. Finally, when applying diffusion models in the latent, we propose a simplified forward diffusion process, which adds Gaussian noises of linear standard deviation with respect to time. We demonstrate through theory and experiments that this approach can reduce the errors in the reverse process, thus improving sampling speed.
@@
Recognizing the absence of unified and comprehensive evaluations for tabular data synthesis methods, we perform extensive experiments, which involve comparing TabSyn with seven state-of-the-art methods on six mixed-type tabular datasets using five distinct evaluation metrics. The experimental results demonstrate that TabSyn consistently outperforms previous methods (see Figure 1). Specifically, TabSyn reduces the average errors in column-wise distribution shape estimation (i.e., single density) and pair-wise column correlation estimation (i.e., pair correlation) tasks by $86\%$ and $67\%$ than the most competitive baselines, demonstrating the superiority of TabSyn. Furthermore, we demonstrate that TabSyn achieves competitive performance across two downstream tabular data tasks, machine learning efficiency and missing value imputation. Specifically, the well-learned unconditional TabSyn is able to be applied to missing value imputation without retraining. Moreover, thorough ablation studies and visualization case studies substantiate the rationale and effectiveness of our developed approach.
@@ 2 Related Works
Generative models for tabular data have become increasingly important and have widespread applications Assefa et al. (2021); Zheng & Charoenphakdee (2022); Hernandez et al. (2022). To deal with the imbalanced categorical features in tabular data, Xu et al. (2019) proposes CTGAN and TVAE based on the popular Generative Adversarial Networks (Goodfellow et al., 2014) and VAE (Kingma & Welling, 2013), respectively. Multiple advanced methods have been proposed for synthetic tabular data generation in the past year. Specifically, GOGGLE (Liu et al., 2023b) became the first to explicitly model the dependency relationship between columns, proposing a VAE-based model using graph neural networks as the encoder and decoder models. Inspired by the success of large language models in modeling the distribution of natural languages, GReaT transformed each row in a table into a natural sentence and learned sentence-level distributions using auto-regressive GPT2. STaSy (Kim et al., 2023), TabDDPM (Kotelnikov et al., 2023), and CoDi (Lee et al., 2023) concurrently applied the popular diffusion-based generative models for synthetic tabular data generation.
@@
While generative models in the data space have achieved significant success, latent generative models have demonstrated several advantages, including more compact and disentangled representations, robustness to noise and greater flexibility in controlling generated styles (van den Oord et al., 2017; Razavi et al., 2019; Esser et al., 2021). Recently, the Latent Diffusion Models (LDM) (Rombach et al., 2022; Vahdat et al., 2021) have achieved great success in image generation as they exhibit better scaling properties and expressivity than the vanilla diffusion models in the data space (Ho et al., 2020; Song et al., 2021b; Karras et al., 2022). The success of LDMs in image generation has also inspired their applications in video (Blattmann et al., 2023) and audio data (Liu et al., 2023a). To the best of our knowledge, the proposed work is the first to explore the application of latent diffusion models for general tabular data generation tasks.
@@ 3.1 Problem Definition of Tabular Data Generation
Let $M_{{\rm num}}$ and $M_{{\rm cat}}$ ($M=M_{{\rm num}}+M_{{\rm cat}}$) be the number of numerical columns and categorical columns, respectively. Each row is represented as a vector of numerical features and categorical features ${\bm{x}}=[{\bm{x}}^{{\rm num}},{\bm{x}}^{{\rm cat}}]$, where ${\bm{x}}^{{\rm num}}\in{\mathbb{R}}^{M_{{\rm num}}}$ and ${\bm{x}}^{{\rm cat}}\in{\mathbb{R}}^{M_{{\rm cat}}}$. Specifically, the $i$-th categorical attribute has $C_{i}$ finite candidate values, therefore we have $x_{i}^{{\rm cat}}\in\{1,\cdots,{C_{i}}\},\forall i$.
This paper focuses on the unconditional generation task. With a tabular dataset ${\mathcal{T}}=\{{\bm{x}}\}$, we aim to learn a parameterized generative model $p_{\theta}({\mathcal{T}})$, with which realistic and diverse synthetic tabular data $\hat{{\bm{x}}}\in\hat{{\mathcal{T}}}$ can be generated.
@@ 3.2 AutoEncoding for Tabular Data
Tabular data is highly structured of mixed-type column features, with columns having specific meanings, and different columns are highly dependent on each other. These characteristics make it challenging to design an approximate encoder to model and effectively utilize the rich relationships between columns. Motivated by the successes of applying Transformers for classification/regression of tabular data (Gorishniy et al., 2021), we first learn a unique tokenizer for each column, and then the token(column)-wise representations are fed into a Transformer for capturing the intricate relationships among columns.
@@ 3.3 Score-based Generative Modeling in the Latent Space
$\mathcal{L}={\mathbb{E}}_{{\bm{z}}_{0}\sim p({\bm{z}}_{0})}{\mathbb{E}}_{t\sim p(t)}{\mathbb{E}}_{\bm{\varepsilon}\sim{\mathcal{N}}(\bm{0},{\bm{I}})}\|\bm{\epsilon}_{\theta}({\bm{z}}_{t},t)-\bm{\varepsilon})\|_{2}^{2},\;~{}\text{where}~{}{\bm{z}}_{t}={\bm{z}}_{0}+\sigma(t)\bm{\varepsilon},$ (7)
@@ 4 Benchmarking Synthetic Tabular Data Generation Algorithms
In this section, we conduct experiments to evaluate TabSyn against alternative methods from various data quality measurements. We deliver a unified and comprehensive comparison environment for tabular data synthesis methods and have released the code base [^1] https://github.com/amazon-science/tabsyn.
@@ 4.1 Experimental Setups
Datasets. We select six real-world tables, including numerical and categorical column attributes: Adult, Default, Shoppers, Magic, Faults, Beijing, and News. Table 6 provides the overall statistics of these datasets, and the detailed descriptions can be found in Appendix E.1.
Baselines. We compare the proposed TabSyn with seven existing synthetic tabular data generation methods. The first two are classical GAN and VAE models: CTGAN (Xu et al., 2019) and TVAE (Xu et al., 2019). Additionally, we evaluate five SOTA methods introduced recently: GOGGLE (Liu et al., 2023b), a VAE-based method; GReaT (Borisov et al., 2023), a language model variant; and three diffusion-based methods: STaSy (Kim et al., 2023), TabDDPM (Kotelnikov et al., 2023), and CoDi (Lee et al., 2023). Notably, these approaches were nearly simultaneously introduced, limiting opportunities for extensive comparison. Our paper fills this gap by providing the first comprehensive evaluation of their performance in a standardized setting.
@@
Implementation details. The reported results are averaged over $20$ randomly sampled synthetic data. The implementation details are in Appendix E.
@@ 4.2 Estimating Low-order Statistics of Data Density
We start by evaluating the synthetic data’s capacity to estimate the density of single columns and the correlation between pairs of columns.
@@
Table 1: Error rate (%) of column-wise density estimation. Bold Face represents the best score on each dataset. Lower values indicate more accurate estimation (superior results). TabSyn outperforms the best baseline model by $86.0\%$ on average.
Method Adult Default Shoppers Magic Beijing News Average Ranking
CTGAN $16.84$$\pm$ $0.03$ $16.83$$\pm$$0.04$ $21.15$±$\pm 0.10$ $9.81$±$\pm 0.08$ $21.39$±$\pm 0.05$ $16.09$±$\pm 0.02$ $17.02$ $7$
TVAE $14.22$±$\pm 0.08$ $10.17$$\pm$$0.05$ $24.51$±$\pm 0.06$ $8.25$±$\pm 0.06$ $19.16$±$\pm 0.06$ $16.62$±$\pm 0.03$ $15.49$ $5$
[…]
STaSy $11.29$±$\pm 0.06$ $5.77$±$\pm 0.06$ $9.37$±$\pm 0.09$ $6.29$±$\pm 0.13$ $6.71$±$\pm 0.03$ $6.89$±$\pm 0.03$ $7.72$ $2$
CoDi $21.38$±$\pm 0.06$ $15.77$$\pm$ $0.07$ $31.84$±$\pm 0.05$ $11.56$±$\pm 0.26$ $16.94$±$\pm 0.02$ $32.27$±$\pm 0.04$ $21.63$ $8$
[…]
TabSyn $\bf{{\color[rgb]{0.8,0.25,0.33}0.58}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.06}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}0.85}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.04}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.43}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.24}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}0.88}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.09}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.12}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.05}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.64}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.04}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.08}}$ $1$
@@
Method Adult Default Shoppers Magic Beijing News Average Ranking
CTGAN $20.23$±$\pm 1.20$ $26.95$±$\pm 0.93$ $13.08$±$\pm 0.16$ $7.00$±$\pm 0.19$ $22.95$±$\pm 0.08$ $5.37$±$\pm 0.05$ $15.93$ $5$
TVAE $14.15$±$\pm 0.88$ $19.50$$\pm$$0.95$ $18.67$±$\pm 0.38$ $5.82$±$\pm 0.49$ $18.01$±$\pm 0.08$ $6.17$±$\pm 0.09$ $13.72$ $4$
GOGGLE $45.29$ $21.94$ $23.90$ $9.47$ $45.94$ $23.19$ $28.28$ $7$
GReaT $17.59$±$\pm 0.22$ $70.02$$\pm$$0.12$ $45.16$±$\pm 0.18$ $10.23$±$\pm 0.40$ $59.60$±$\pm 0.55$ $44.24$ $8$
STaSy $14.51$±$\pm 0.25$ $5.96$$\pm$$0.26$ $8.49$±$\pm 0.15$ $6.61$±$\pm 0.53$ $8.00$±$\pm 0.10$ $3.07$±$\pm 0.04$ $7.77$ $3$
CoDi $22.49$±$\pm 0.08$ $68.41$$\pm$$0.05$ $17.78$±$\pm 0.11$ $6.53$±$\pm 0.25$ $7.07$±$\pm 0.15$ $11.10$±$\pm 0.01$ $22.23$ $6$
TabDDPM $3.01$±$\pm 0.25$ $4.89$±$\pm 0.10$ $6.61$±$\pm 0.16$ $1.70$±$\pm 0.22$ $2.71$±$\pm 0.09$ $13.16$±$\pm 0.11$ $5.34$ $2$
TabSyn $\bf{{\color[rgb]{0.8,0.25,0.33}1.54}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.27}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}2.05}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.12}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}2.07}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.21}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.06}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.31}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}2.24}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.28}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.44}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.03}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.73}}$ $1$

### v3

@@ Abstract
Recent advances in tabular data generation have greatly enhanced synthetic data quality. However, extending diffusion models to tabular data is challenging due to the intricately varied distributions and a blend of data types of tabular data. This paper introduces TabSyn, a methodology that synthesizes tabular data by leveraging a diffusion model within a variational autoencoder (VAE) crafted latent space. The key advantages of the proposed TabSyn include (1) Generality: the ability to handle a broad spectrum of data types by converting them into a single unified space and explicitly capturing inter-column relations; (2) Quality: optimizing the distribution of latent embeddings to enhance the subsequent training of diffusion models, which helps generate high-quality synthetic data, (3) Speed: much fewer number of reverse steps and faster synthesis speed than existing diffusion-based methods. Extensive experiments on six datasets with five metrics demonstrate that TabSyn outperforms existing methods. Specifically, it reduces the error rates by 86% and 67% for column-wise distribution and pair-wise column correlation estimations compared with the most competitive baselines. The code has been made available at https://github.com/amazon-science/tabsyn.
@@ 1 Introduction
Figure 1: Our TabSyn consistently outperforms SOTA tabular data generation methods across five data quality metrics.
@@
This paper presents TabSyn, a principled approach for tabular data synthesis. To handle mixed-typed inputs, TabSyn first transforms raw tabular data into a continuous embedding space, where well-developed diffusion models with Gaussian noises become feasible. Subsequently, we learn a score-based diffusion model in the embedding space to capture the distribution of latent embeddings. To learn an informative, smoothed latent space while maintaining the decoder’s reconstruction ability, we specifically designed a Variational AutoEncoder (VAE (Kingma & Welling, 2013)) model for tabular-structured data. Our proposed VAE model includes 1) Transformer-architecture encoders and decoders for modeling inter-column relationships and obtaining token-level representations, facilitating token-level tasks. 2) Adaptive loss weighting to dynamically adjust the reconstruction loss weights and KL-divergence weights, allowing the model to improve reconstruction performance gradually while maintaining a regularized embedding space. 3) Finally, when applying diffusion models in the latent space, we adopt a simplified forward diffusion process, which adds Gaussian noises of linear standard deviation with respect to time. We demonstrate through theoretical analysis and empirical justifications that this approach can reduce the errors in the reverse process, thus improving sampling speed.
@@
Recognizing the absence of unified and comprehensive evaluations for synthetic tabular data (Du & Li, 2024), we perform extensive experiments, which involve comparing TabSyn with seven state-of-the-art methods on six mixed-type tabular datasets using over five distinct evaluation metrics. The experimental results demonstrate that TabSyn consistently outperforms previous methods (see Figure 1). Specifically, TabSyn reduces the average errors in column-wise distribution shape estimation (i.e., single density) and pair-wise column correlation estimation (i.e., pair correlation) tasks by $86\%$ and $67\%$ than the most competitive baselines. Furthermore, we demonstrate that TabSyn achieves competitive performance across two downstream tabular data tasks, machine learning efficiency and missing value imputation. Specifically, the well-learned unconditional TabSyn is able to be applied to missing value imputation without retraining. Moreover, thorough ablation studies and visualization case studies substantiate the rationale and effectiveness of our developed approach.
@@ 2 Related Works
Generative models for tabular data have become increasingly important and have widespread applications Assefa et al. (2021); Zheng & Charoenphakdee (2022); Hernandez et al. (2022). To deal with the imbalanced categorical features, Xu et al. (2019) proposes CTGAN and TVAE based on the popular Generative Adversarial Networks (Goodfellow et al., 2014) and VAE (Kingma & Welling, 2013), respectively. Multiple advanced methods have been proposed for synthetic tabular data generation in the past year. Specifically, GOGGLE (Liu et al., 2023b) became the first to explicitly model the dependency relationship between columns, proposing a VAE-based model using graph neural networks as the encoder and decoder models. Inspired by the success of large language models in modeling the distribution of natural languages, GReaT transformed each row in a table into a natural sentence and learned sentence-level distributions using auto-regressive GPT2. In recent years, the physical diffusion process has inspired a lot of advanced research in deep learning. For example, DIFFormer (Wu et al., 2023) develops a scalable Transformer model for geometric data via a constrained diffusion process, and the Denoising Diffusion models have achieved great success in image generation (Ho et al., 2020). STaSy (Kim et al., 2023), TabDDPM (Kotelnikov et al., 2023), and CoDi (Lee et al., 2023) concurrently applied the popular diffusion-based generative models for synthetic tabular data generation.
@@
While generative models in the data space have achieved significant success, latent generative models have demonstrated several advantages, including more compact and disentangled representations, robustness to noise, and greater flexibility in controlling generated styles (van den Oord et al., 2017; Razavi et al., 2019; Esser et al., 2021). For example, the recent GAN literature (Li et al., 2022) has demonstrated superior controllability via adversarial learning in the latent space. Recently, the Latent Diffusion Models (LDM) (Rombach et al., 2022; Vahdat et al., 2021) have achieved great success in image generation as they exhibit better scaling properties and expressivity than the vanilla diffusion models in the data space (Ho et al., 2020; Song et al., 2021b; Karras et al., 2022). The success of LDMs in image generation has also inspired their applications in video (Blattmann et al., 2023) and audio data (Liu et al., 2023a). To the best of our knowledge, the proposed work is the first to explore the application of latent diffusion models for general tabular data generation tasks.
@@ 3.1 Problem Definition of Tabular Data Generation
Let $M_{{\rm num}}$ and $M_{{\rm cat}}$ be the number of numerical columns and categorical columns, respectively. Each row is represented as a vector of numerical features and categorical features ${\bm{x}}=[{\bm{x}}^{{\rm num}},{\bm{x}}^{{\rm cat}}]$, where ${\bm{x}}^{{\rm num}}\in{\mathbb{R}}^{M_{{\rm num}}}$ and ${\bm{x}}^{{\rm cat}}\in{\mathbb{R}}^{M_{{\rm cat}}}$. Specifically, the $i$-th categorical attribute has $C_{i}$ finite candidate values, therefore we have $x_{i}^{{\rm cat}}\in\{1,\cdots,{C_{i}}\},\forall i$. This paper focuses on the unconditional generation task. With a tabular dataset ${\mathcal{T}}=\{{\bm{x}}\}$, we aim to learn a parameterized generative model $p_{\theta}({\mathcal{T}})$, with which realistic and diverse synthetic tabular data $\hat{{\bm{x}}}\in\hat{{\mathcal{T}}}$ can be generated.
@@ 3.2 AutoEncoding for Tabular Data
Tabular data is highly structured of mixed-type column features, with different columns having distinct meanings and being highly dependent on each other. These characteristics make it challenging to design an approximate encoder to model and effectively utilize the rich relationships between columns. Motivated by the successes of Transformers in classification/regression of tabular data (Gorishniy et al., 2021), we first learn a unique tokenizer for each column, and then the token(column)-wise representations are fed into a Transformer for capturing the intricate relationships among columns.
@@ 3.3 Score-based Generative Modeling in the Latent Space
$\mathcal{L}={\mathbb{E}}_{{\bm{z}}_{0}\sim p({\bm{z}}_{0})}{\mathbb{E}}_{t\sim p(t)}{\mathbb{E}}_{\bm{\varepsilon}\sim{\mathcal{N}}(\bm{0},{\bm{I}})}\|\bm{\epsilon}_{\theta}({\bm{z}}_{t},t)-\bm{\varepsilon}\|_{2}^{2},\;~{}\text{where}~{}{\bm{z}}_{t}={\bm{z}}_{0}+\sigma(t)\bm{\varepsilon},$ (7)
@@ 4.1 Experimental Setups
Datasets. We select six real-world tabular datasets consisting of both numerical and categorical attributes: Adult, Default, Shoppers, Magic, Faults, Beijing, and News. Table 6 provides the overall statistics of these datasets, and the detailed descriptions can be found in Appendix E.1.
Baselines. We compare the proposed TabSyn with seven existing synthetic tabular data generation methods. The first two are classical GAN and VAE models: CTGAN (Xu et al., 2019) and TVAE (Xu et al., 2019). Additionally, we evaluate five SOTA methods introduced recently: GOGGLE (Liu et al., 2023b), a VAE-based method; GReaT (Borisov et al., 2023), a language model variant; and three diffusion-based methods: STaSy (Kim et al., 2023), TabDDPM (Kotelnikov et al., 2023), and CoDi (Lee et al., 2023). Notably, these approaches were nearly simultaneously introduced, limiting opportunities for extensive comparison. For reference, we also compare with the representative interpolation-based method SMOTE (Chawla et al., 2002). Our paper fills this gap by providing the first comprehensive evaluation of their performance in a standardized setting.
@@ 4.2 Estimating Low-order Statistics of Data Density
Table 1: Error rate (%) of column-wise density estimation. Bold Face represents the best score on each dataset. Lower values indicate more accurate estimation (superior results). TabSyn outperforms the best generative baseline model by $86.0\%$ on average.
Method Adult Default Shoppers Magic Beijing News Average
SMOTE $1.60$±$\pm 0.23$ $1.48$±$\pm 0.15$ $2.68$±$\pm 0.19$ $0.91$±$\pm 0.05$ $1.85$±$\pm 0.21$ $5.31$±$\pm 0.46$ $2.30$
CTGAN $16.84$$\pm$ $0.03$ $16.83$$\pm$$0.04$ $21.15$±$\pm 0.10$ $9.81$±$\pm 0.08$ $21.39$±$\pm 0.05$ $16.09$±$\pm 0.02$ $17.02$
TVAE $14.22$±$\pm 0.08$ $10.17$$\pm$$0.05$ $24.51$±$\pm 0.06$ $8.25$±$\pm 0.06$ $19.16$±$\pm 0.06$ $16.62$±$\pm 0.03$ $15.49$
GOGGLE1 $16.97$ $17.02$ $22.33$ $1.90$ $16.93$ $25.32$ $16.74$
GReaT2 $12.12$$\pm$$0.04$ $19.94$$\pm$$0.06$ $14.51$±$\pm 0.12$ $16.16$±$\pm 0.09$ $8.25$±$\pm 0.12$ OOM $14.20$
STaSy $11.29$±$\pm 0.06$ $5.77$±$\pm 0.06$ $9.37$±$\pm 0.09$ $6.29$±$\pm 0.13$ $6.71$±$\pm 0.03$ $6.89$±$\pm 0.03$ $7.72$
CoDi $21.38$±$\pm 0.06$ $15.77$$\pm$ $0.07$ $31.84$±$\pm 0.05$ $11.56$±$\pm 0.26$ $16.94$±$\pm 0.02$ $32.27$±$\pm 0.04$ $21.63$
TabDDPM3 $1.75$±$\pm 0.03$ $1.57$$\pm$ $0.08$ $2.72$±$\pm 0.13$ $1.01$±$\pm 0.09$ $1.30$±$\pm 0.03$ $78.75$±$\pm 0.01$ $14.52$
TabSyn $\bf{{\color[rgb]{0.8,0.25,0.33}0.58}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.06}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}0.85}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.04}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.43}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.24}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}0.88}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.09}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.12}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.05}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.64}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.04}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.08}}$
@@
Method Adult Default Shoppers Magic Beijing News Average
SMOTE $3.28$±$\pm 0.29$ $8.41$±$\pm 0.38$ $3.56$±$\pm 0.22$ $3.16$±$\pm 0.41$ $2.39$±$\pm 0.35$ $5.38$±$\pm 0.76$ $4.36$
CTGAN $20.23$±$\pm 1.20$ $26.95$±$\pm 0.93$ $13.08$±$\pm 0.16$ $7.00$±$\pm 0.19$ $22.95$±$\pm 0.08$ $5.37$±$\pm 0.05$ $15.93$
TVAE $14.15$±$\pm 0.88$ $19.50$$\pm$$0.95$ $18.67$±$\pm 0.38$ $5.82$±$\pm 0.49$ $18.01$±$\pm 0.08$ $6.17$±$\pm 0.09$ $13.72$
GOGGLE $45.29$ $21.94$ $23.90$ $9.47$ $45.94$ $23.19$ $28.28$
GReaT $17.59$±$\pm 0.22$ $70.02$$\pm$$0.12$ $45.16$±$\pm 0.18$ $10.23$±$\pm 0.40$ $59.60$±$\pm 0.55$ OOM $44.24$
STaSy $14.51$±$\pm 0.25$ $5.96$$\pm$$0.26$ $8.49$±$\pm 0.15$ $6.61$±$\pm 0.53$ $8.00$±$\pm 0.10$ $3.07$±$\pm 0.04$ $7.77$
CoDi $22.49$±$\pm 0.08$ $68.41$$\pm$$0.05$ $17.78$±$\pm 0.11$ $6.53$±$\pm 0.25$ $7.07$±$\pm 0.15$ $11.10$±$\pm 0.01$ $22.23$
TabDDPM $3.01$±$\pm 0.25$ $4.89$±$\pm 0.10$ $6.61$±$\pm 0.16$ $1.70$±$\pm 0.22$ $2.71$±$\pm 0.09$ $13.16$±$\pm 0.11$ $5.34$
TabSyn $\bf{{\color[rgb]{0.8,0.25,0.33}1.54}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.27}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}2.05}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.12}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}2.07}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.21}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.06}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.31}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}2.24}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.28}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.44}}$±$\bf{{\color[rgb]{0.8,0.25,0.33}\pm 0.03}}$ $\bf{{\color[rgb]{0.8,0.25,0.33}1.73}}$